OpenStudy (anonymous):

Find all solutions to the equation in the interval [0, 2π). cos 2x - cos x = 0 0, 2pi/3, 4pi/3 pi/6, 5pi/6, 3pi/2 0, pi/2, 7pi/6, 11pi/6 No solution

OpenStudy (ranga):

Is the first term cos(two x) or cos-squared-x ?

OpenStudy (anonymous):

cos 2x

OpenStudy (ranga):

Use the trig identity: cos(2x) = 2cos^2(x) - 1 Then it will be a quadratic equation in cos which can be solved

OpenStudy (anonymous):

i don't know how to do tht

OpenStudy (ranga):

cos(2x) - cos(x) = 0 2cos^2(x) - 1 - cos(x) = 0 2cos^2(x) - cos(x) - 1= 0 Let u = cos(x) 2u^2 - u - 1 = 0 2u^2 - 2u + u - 1 = 0 2u(u-1) + 1(u-1) = 0 (2u+1)(u-1) = 0 u = -1/2 or u = 1 put back cos(x) in place of u: cos(x) = -1/2 or cos(x) = 1 Find out all x in the interval [0,2pi] where cos(x) = -1/2 or 1

OpenStudy (anonymous):

i don't understand what to do after that.....

OpenStudy (anonymous):

@ranga

OpenStudy (ranga):

Find out all x in the interval [0,2pi] where cos(x) = -1/2 or 1 Look up the unit circle. cos(x) is -1/2 when x = 2pi/3, 4pi/3 cos(x) is 1 when x = 0. So x = 0, 2pi/3, 4pi/3